# Existence and uniqueness of the canonical Brownian motion in non-simple conformal loop ensemble gaskets

### v1

@@ 1.2. Main results
Throughout the paper we let $\kappa^{\prime}\in(4,8)$ be fixed. We let
@@
    There exists a constant $\alpha>0$ so that the following holds. Let $U\subseteq\mathbf{C}$ be open, simply connected, and $\lambda>0$. Then
    $(\mathcal{E}^{V}(\cdot,\cdot;\Gamma_{U^{*}}))_{V\in\mathfrak{C}_{U}}=(\lambda^{\alpha}\mathcal{E}^{\lambda V}(S_{\lambda}\cdot,S_{\lambda}\cdot;\lambda\Gamma_{U^{*}}))_{V\in\mathfrak{C}_{U}}\quad\text{a.s.,}$
@@
For each $\kappa^{\prime}\in(4,8)$ there exists a unique (up to a deterministic factor) ${\rm CLE}_{\kappa^{\prime}}$ resistance form in the sense of Definition 1.1. The exponent $\alpha$ depends only on $\kappa^{\prime}$ and satisfies $\alpha\in[{d_{{\mathrm{dbl}}}},{d_{{\rm SLE}}}]$.
@@
      There exists a constant $\alpha>0$ so that the following holds. For each open, simply connected $U\subseteq\mathbf{C}$ and $\lambda>0$, the functions $F_{U}$, $F_{\lambda U}$ can be chosen so that $F_{\lambda U}=(S_{\lambda})_{*}F_{U}(\lambda^{-1}\cdot)$ where $S_{\lambda}X(t)=\lambda X(\lambda^{-d_{\rm CLE}-\alpha}t)$.
@@
By [Kig12, Theorem 10.4], for each compact resistance metric space and each finite Borel measure with full support the associated Hunt process possesses a continuous heat kernel. Moreover, by [Kig12, (10.4)] and the regularity estimates for the resistance metric and the measure (Lemma 4.5 and Proposition 2.7), we have the on-diagonal upper bound $p(t,x,x)\leq t^{-d_{\rm CLE}/(d_{\rm CLE}+\alpha)-o(1)}$ as $t\searrow 0$ where $\alpha$ is the exponent in Theorem 1.2. We expect that the matching lower bound also holds, i.e. the spectral dimension of the ${\rm CLE}_{\kappa^{\prime}}$ Brownian motion is $2d_{\rm CLE}/(d_{\rm CLE}+\alpha)$.
@@
    There exists a constant $\alpha>0$ so that for each $\lambda>0$, we have
    $\mathcal{E}(\cdot,\cdot;\Gamma^{\mathbf{C}})=\lambda^{\alpha}\mathcal{E}(S_{\lambda}\cdot,S_{\lambda}\cdot;\lambda\Gamma^{\mathbf{C}})\quad\text{a.s.,}$
@@
    There exists a constant $\alpha>0$ so that the following holds. For each simply connected domain $D\subseteq\mathbf{C}$ and $\lambda>0$, we have
    $\mathcal{E}(\cdot,\cdot;\Gamma^{D})=\lambda^{\alpha}\mathcal{E}(S_{\lambda}\cdot,S_{\lambda}\cdot;\lambda\Gamma^{D})\quad\text{a.s.,}$
@@
The equivalence of Theorems 1.5 and 1.6 with Theorem 1.2 resp. 1.4 follows by the exact same argument as in [MY25, Theorems 1.7 and 1.8].
@@ 1.4. Outline
In Appendix A we establish some strong concentration and homogeneity results for the ${\rm CLE}_{\kappa^{\prime}}$ gasket measure (focusing on the case $\kappa^{\prime}\in(4,8)$). This will serve as an input into the proof of tightness in Section 3 and is of independent interest.
@@
We assume throughout the paper that $\kappa^{\prime}\in(4,8)$ is fixed, and we write $\kappa=16/\kappa\in(2,4)$. We always consider nested ${\rm CLE}_{\kappa^{\prime}}$ in the proofs (which is required to use the resampling tools from [AMY25a]; see Section 2.2.2).
@@ 2.2.3. Resampling target pivotals
We will now describe the results of [AMY25a, Section 5], which is focused on using a “resampling procedure” that alters the connections in a (multichordal) ${\rm CLE}_{\kappa}$. Let us now describe more precisely what we mean by a resampling procedure. Suppose that $\Gamma$ is a multichordal ${\rm CLE}_{\kappa}$ in $(D;\underline{x};\beta)$, and let $W\subseteq D$. Then we can construct another multichordal ${\rm CLE}_{\kappa}$ in $(D;\underline{x};\beta)$ coupled with $\Gamma$ by repeatedly applying the following procedure. We select $(U,V)\in{\mathfrak{P}}_{D}$ with $V\subseteq W$ randomly in a way that is independent of the CLE configuration within $W$, and let $\widetilde{\Gamma}$ be such that $\widetilde{\Gamma}_{\mathrm{out}}^{*,V,U}=\Gamma_{\mathrm{out}}^{*,V,U}$, and $\widetilde{\Gamma}_{\mathrm{in}}^{*,V,U}$ is sampled from its conditional law given $\widetilde{\Gamma}_{\mathrm{out}}^{*,V,U}$ (independently of $\Gamma_{\mathrm{in}}^{*,V,U}$). We say that $\Gamma^{\mathrm{res}}$ is a resampling of $\Gamma$ within $W$ if it is obtained by applying such a procedure a (random) number of times.
@@ 2.3. CLE gasket measure
Let $D\subseteq\mathbf{C}$ be a simply connected domain, let $\Gamma$ be a non-nested ${\rm CLE}_{\kappa^{\prime}}$ in $D$, and let $\Upsilon_{\Gamma}$ be its gasket. Let $d_{\rm CLE}$ be as in (1.1). It is shown in [MS22] that there exists a unique (up to a deterministic constant) measure $\mu(\cdot;\Gamma)$ on $\Upsilon_{\Gamma}$ satisfying the following properties.
  • We have $\mathbf{E}[\mu(K;\Gamma)]<\infty$ for each compact set $K\subseteq D$.
  • Let $U\subseteq D$ be open, simply connected, and let $U^{*}\subseteq U$ be the set of points that are not on or inside a loop $\mathcal{L}\in\Gamma$ with $\mathcal{L}\cap D\setminus U\neq\emptyset$. For each connected component $V$ of $U^{*}$, let $\Gamma_{V}\subseteq\Gamma$ be the loops contained in $\overline{V}$. Then the measure $\mu(\cdot\cap V;\Gamma)$ is determined by $\Gamma_{V}$ and almost surely
    $\mu(\varphi(A);\Gamma_{V})=\int_{A}\lvert\varphi^{\prime}(z)\rvert^{d_{\rm CLE}}\mu(dz;\varphi^{-1}(\Gamma_{V}))$
    where $\varphi\colon D\to V$ is a conformal transformation.
We state the following regularity properties of the ${\rm CLE}_{\kappa^{\prime}}$ measure whose proofs will be given in Appendix A (and have not appeared in other works).
Proposition 2.6.
Let $\Gamma$ be a ${\rm CLE}_{\kappa^{\prime}}$ in a simply connected domain $D\subseteq\mathbf{C}$ and let $\mu(\cdot;\Gamma)$ denote the gasket measure on $\Upsilon_{\Gamma}$. Fix any $a>0$ and a compact set $K\subseteq D$. For each $r>0$, let $E_{r}$ be the event that
$\sup_{z\in K}\mu(B_{E}(z,r);\Gamma)\leq r^{d_{\rm CLE}-a}.$
Then $\mathbf{P}[E_{r}^{c}]=o^{\infty}(r)$ as $r\searrow 0$.
Proposition 2.7.
Let $\Gamma$ be a ${\rm CLE}_{\kappa^{\prime}}$ in a simply connected domain $D\subseteq\mathbf{C}$ and let $\mu(\cdot;\Gamma)$ denote the gasket measure on $\Upsilon_{\Gamma}$. Fix any $a>0$ and a compact set $K\subseteq D$. For each $r>0$, let $E_{r}$ be the event that the following holds. Let $x\in\Upsilon_{\Gamma}\cap K$ and $w\in\partial B_{\mathrm{path}}(x,r)$. Let $B_{x,w}\subseteq\overline{B_{E}}(x,r)$ be a simply connected set containing all simple admissible paths within $\overline{B_{E}}(x,r)$ from $x$ to $w$. Then
$\mu(B_{x,w};\Gamma)\geq r^{d_{\rm CLE}+a}\quad\text{for each such }x,w.$
Then $\mathbf{P}[E_{r}^{c}]=o^{\infty}(r)$ as $r\searrow 0$.
@@ 3.3.1. Approximation scheme
Let $\mu(\cdot;\Gamma)$ denote the gasket measure (see Section 2.3), and let $\mathfrak{V}_{\epsilon}$ be a Poisson point process with intensity measure $\lambda_{\epsilon}\mu(\cdot;\Gamma)$. As we will explain just below, the concentration results from Appendix A for $\mu(\cdot;\Gamma)$ imply that every Euclidean ball of radius $\epsilon$ is likely to contain many points of $\mathfrak{V}_{\epsilon}$ (i.e., at least a negative power of $\epsilon$).
@@ 4.3. Good annulus event
If we condition on the chords of a multichordal ${\rm CLE}_{\kappa^{\prime}}$, the remainder has the conditional law of an ordinary ${\rm CLE}_{\kappa}$. Hence, there exists some $\widetilde{D}$ with $\operatorname{dist_{E}}(z,\partial\widetilde{D})\in[2^{-j},2^{-j+1}]$ such that
@@ 5. Uniqueness
In the previous section, we proved that if we have two weak ${\rm CLE}_{\kappa^{\prime}}$ resistance forms $(\mathcal{E}^{V}(\cdot,\cdot;\Gamma))_{V\in\mathfrak{C}}$ and $(\widetilde{\mathcal{E}}^{V}(\cdot,\cdot;\Gamma))_{V\in\mathfrak{C}}$ with comparable scaling factors, then they are bi-Lipschitz equivalent with deterministic constants $0<c_{1}\leq c_{2}<\infty$. We show in Section 5.1 that in this setting one can take $c_{1}=c_{2}$. Then we show in Section 5.2 that there is a unique scaling exponent $\alpha>0$ depending only on $\kappa^{\prime}$ such that every ${\rm CLE}_{\kappa^{\prime}}$ resistance form satisfies scaling covariance with this exponent $\alpha>0$. This in particular lifts the assumption (4.1) on the comparability of the scaling factors.
@@ 5.2. Uniqueness of the exponent
In this section we show the scaling covariance of ${\rm CLE}_{\kappa^{\prime}}$ resistance forms with a scaling factor $\lambda^{\alpha}$ where $\alpha>0$ depends only on $\kappa^{\prime}$ (Proposition 5.5).
@@
There exists $\alpha\in[{d_{{\mathrm{dbl}}}},{d_{{\rm SLE}}}]$ so that the following is true. Suppose that $(\mathcal{E}^{V}(\cdot,\cdot;\Gamma))_{V\in\mathfrak{C}}$ is a weak ${\rm CLE}_{\kappa^{\prime}}$ resistance form, then it is a (strong) ${\rm CLE}_{\kappa^{\prime}}$ resistance form with the scaling exponent $\alpha$.
We will prove Proposition 5.5 in two steps. The first step, stated and proved in Lemma 5.6 just below, is to show that each weak ${\rm CLE}_{\kappa^{\prime}}$ resistance form is a (strong) ${\rm CLE}_{\kappa^{\prime}}$ resistance form with some scaling exponent $\alpha>0$. We will then prove that the value of $\alpha>0$ does not depend on the choice of the ${\rm CLE}_{\kappa^{\prime}}$ resistance form, but only on the choice of $\kappa^{\prime}\in(4,8)$.
@@
Suppose that we have the setup of Section 4.1 but condition on the event that $\mathcal{L}\cap\partial\mathbf{D}=\emptyset$. [^7] This conditioning implies that for each $\lambda\in(0,1]$, the law of $\lambda\Gamma$ is absolutely continuous with respect to the law of $\Gamma$, see [MY25, Lemma 7.2]. Let $(\mathcal{E}^{V}(\cdot,\cdot;\Gamma))_{V\in\mathfrak{C}}$ be a weak ${\rm CLE}_{\kappa^{\prime}}$ resistance form. Then there exists $\alpha\in[{d_{{\mathrm{dbl}}}},{d_{{\rm SLE}}}]$ such that the following holds. Let $\lambda\in(0,1]$ and define $(\widetilde{\mathcal{E}}^{V}(f,f;\Gamma))_{V\in\mathfrak{C}}=(\mathcal{E}^{\lambda V}(f(\lambda^{-1}\cdot),f(\lambda^{-1}\cdot);\lambda\Gamma))_{V\in\mathfrak{C}}$. Then almost surely $\widetilde{\mathcal{E}}^{V}(\cdot,\cdot;\Gamma)=\lambda^{-\alpha}\mathcal{E}^{V}(\cdot,\cdot;\Gamma)$ for each $V\in\mathfrak{C}$.
@@
and therefore $c(\lambda^{\prime}\lambda)=c(\lambda^{\prime})c(\lambda)$. Moreover, $\lambda\mapsto c(\lambda)$ is strictly decreasing with $\lim_{\lambda\to 0}c(\lambda)=\infty$ due to Proposition 4.4. It follows that $c(\lambda)=\lambda^{-\alpha}$ for some $\alpha>0$. It also follows from Proposition 4.4 that $\alpha\in[{d_{{\mathrm{dbl}}}},{d_{{\rm SLE}}}]$. ∎
@@
It remains to show that if $\mathcal{E}(\cdot,\cdot;\Gamma)$ and $\widetilde{\mathcal{E}}(\cdot,\cdot;\Gamma)$ are two weak ${\rm CLE}_{\kappa^{\prime}}$ resistance forms with respective scaling exponents $\alpha,\widetilde{\alpha}$, then $\alpha=\widetilde{\alpha}$. Suppose that $\alpha>\widetilde{\alpha}$. Following the proof of Proposition 4.9, there is a covering by good regions $\{U_{l}\}$ analogously to Corollary 4.11 such that
$M^{-1}2^{j\alpha}\leq\frac{\mathcal{E}^{U_{l}}(f,f;\Gamma)}{(\sup_{K_{l}}f-\inf_{K_{l}}f)^{2}}\leq M2^{j\alpha}$
@@
$M^{-1}2^{j\widetilde{\alpha}}\leq\frac{\widetilde{\mathcal{E}}^{U_{l}}(f,f;\Gamma)}{(\sup_{K_{l}}f-\inf_{K_{l}}f)^{2}}\leq M2^{j\widetilde{\alpha}}$

### v2

@@ 1.2. Main results
Throughout the paper we let $\kappa^{\prime}\in(4,8)$ be fixed. We let $d_{\rm CLE}$ be the dimension of the ${\rm CLE}_{\kappa^{\prime}}$ gasket given by (1.1), and
@@
    There exists a constant ${\alpha_{\mathrm{r}}}>0$ so that the following holds. Let $U\subseteq\mathbf{C}$ be open, simply connected, and $\lambda>0$. Then
    $(\mathcal{E}^{V}(\cdot,\cdot;\Gamma_{U^{*}}))_{V\in\mathfrak{C}_{U}}=(\lambda^{{\alpha_{\mathrm{r}}}}\mathcal{E}^{\lambda V}(S_{\lambda}\cdot,S_{\lambda}\cdot;\lambda\Gamma_{U^{*}}))_{V\in\mathfrak{C}_{U}}\quad\text{a.s.,}$
@@
For each $\kappa^{\prime}\in(4,8)$ there exists a unique (up to a deterministic factor) ${\rm CLE}_{\kappa^{\prime}}$ resistance form in the sense of Definition 1.1. The exponent $\alpha_{\mathrm{r}}$} depends only on $\kappa^{\prime}$ and satisfies ${\alpha_{\mathrm{r}}}\in[{d_{{\mathrm{dbl}}}},{d_{{\rm SLE}}}]$.
@@
      There exists a constant ${\alpha_{\mathrm{r}}}>0$ so that the following holds. For each open, simply connected $U\subseteq\mathbf{C}$ and $\lambda>0$, the functions $F_{U}$, $F_{\lambda U}$ can be chosen so that $F_{\lambda U}=(S_{\lambda})_{*}F_{U}(\lambda^{-1}\cdot)$ where $S_{\lambda}X(t)=\lambda X(\lambda^{-d_{\rm CLE}-{\alpha_{\mathrm{r}}}}t)$.
@@
By [Kig12, Theorem 10.4], for each compact resistance metric space and each finite Borel measure with full support the associated Hunt process possesses a continuous heat kernel. Moreover, by [Cro07, Theorem 1], the volume growth $\mu(B_{\mathrm{path}}(x,r);\Gamma)=r^{d_{\rm CLE}+o(1)}$ (see Propositions 2.62.7) and resistance growth ${\mathfrak{R}}(x,y;\Gamma)=d_{\mathrm{path}}(x,y)^{{\alpha_{\mathrm{r}}}+o(1)}$ (see Lemmas 4.54.6) imply the following on-diagonal heat kernel bounds.
Proposition 1.5.
Let $\alpha_{\mathrm{r}}$} be the exponent in Theorem 1.2. Almost surely, the heat kernel of the ${\rm CLE}_{\kappa^{\prime}}$ Brownian motion satisfies
$p(t,x,x)=t^{-d_{\rm CLE}/(d_{\rm CLE}+{\alpha_{\mathrm{r}}})+o(1)}\quad\text{for each }x\in\Upsilon_{\Gamma}\quad\text{as }t\searrow 0.$
That is, the spectral dimension of the ${\rm CLE}_{\kappa^{\prime}}$ Brownian motion is $2d_{\rm CLE}/(d_{\rm CLE}+{\alpha_{\mathrm{r}}})$.
@@
    There exists a constant ${\alpha_{\mathrm{r}}}>0$ so that for each $\lambda>0$, we have
    $\mathcal{E}(\cdot,\cdot;\Gamma^{\mathbf{C}})=\lambda^{{\alpha_{\mathrm{r}}}}\mathcal{E}(S_{\lambda}\cdot,S_{\lambda}\cdot;\lambda\Gamma^{\mathbf{C}})\quad\text{a.s.,}$
@@
    There exists a constant ${\alpha_{\mathrm{r}}}>0$ so that the following holds. For each simply connected domain $D\subseteq\mathbf{C}$ and $\lambda>0$, we have
    $\mathcal{E}(\cdot,\cdot;\Gamma^{D})=\lambda^{{\alpha_{\mathrm{r}}}}\mathcal{E}(S_{\lambda}\cdot,S_{\lambda}\cdot;\lambda\Gamma^{D})\quad\text{a.s.,}$
@@
The equivalence of Theorems 1.6 and 1.7 with Theorem 1.2 resp. 1.4 follows by the exact same argument as in [MY25, Theorems 1.7 and 1.8].
@@ 1.4. Outline
We assume throughout the paper that $\kappa^{\prime}\in(4,8)$ is fixed. We always consider nested ${\rm CLE}_{\kappa^{\prime}}$ in the proofs (which is required to use the resampling tools from [AMY25a]; see Section 2.2.2).
@@ 2.2.3. Resampling target pivotals
We will now describe the results of [AMY25a, Section 5], which is focused on using a “resampling procedure” that alters the connections in a (multichordal) ${\rm CLE}_{\kappa^{\prime}}$. Let us now describe more precisely what we mean by a resampling procedure. Suppose that $\Gamma$ is a multichordal ${\rm CLE}_{\kappa^{\prime}}$ in $(D;\underline{x};\beta)$, and let $W\subseteq D$. Then we can construct another multichordal ${\rm CLE}_{\kappa^{\prime}}$ in $(D;\underline{x};\beta)$ coupled with $\Gamma$ by repeatedly applying the following procedure. We select $(U,V)\in{\mathfrak{P}}_{D}$ with $V\subseteq W$ randomly in a way that is independent of the CLE configuration within $W$, and let $\widetilde{\Gamma}$ be such that $\widetilde{\Gamma}_{\mathrm{out}}^{*,V,U}=\Gamma_{\mathrm{out}}^{*,V,U}$, and $\widetilde{\Gamma}_{\mathrm{in}}^{*,V,U}$ is sampled from its conditional law given $\widetilde{\Gamma}_{\mathrm{out}}^{*,V,U}$ (independently of $\Gamma_{\mathrm{in}}^{*,V,U}$). We say that $\Gamma^{\mathrm{res}}$ is a resampling of $\Gamma$ within $W$ if it is obtained by applying such a procedure a (random) number of times.
@@ 2.3. CLE gasket measure
Let $D\subseteq\mathbf{C}$ be a simply connected domain, let $\Gamma$ be a non-nested ${\rm CLE}_{\kappa^{\prime}}$ in $D$, and let $\Upsilon_{\Gamma}$ be its gasket. Let $d_{\rm CLE}$ be as in (1.1). It is shown in [MS22] that there exists a unique (up to a deterministic constant) measure $\mu(\cdot;\Upsilon_{\Gamma})$ on $\Upsilon_{\Gamma}$ that has finite expectation, is conformally covariant with the exponent $d_{\rm CLE}$, and respects the Markov property of the ${\rm CLE}_{\kappa^{\prime}}$. It is shown later in [MY26] that the measure $\mu(\cdot;\Upsilon_{\Gamma})$ has an explicit description as the Minkowski content measure of $\Upsilon_{\Gamma}$. That is, for a deterministic constant $c>0$, almost surely,
$\mu(A;\Upsilon_{\Gamma})=c\lim_{\delta\searrow 0}\delta^{d_{\rm CLE}-2}\,\operatorname{Leb}{\left(\bigcup_{x\in A\cap\Upsilon_{\Gamma}}B_{E}(x,\delta)\right)}$
for each $A\Subset D$ with, say, rectifiable boundary.
It is further proved that the measure has local dimension $d_{\rm CLE}$ at each point with superpolynomial concentration. In particular, we recall the following upper and lower bounds which will be used later in this work. Suppose that $\Gamma$ is a nested ${\rm CLE}_{\kappa^{\prime}}$ in $D$. Let $(\Upsilon_{k})$ be an enumeration of its gaskets, i.e. each $\Upsilon_{k}$ is the gasket associated with the non-nested ${\rm CLE}_{\kappa^{\prime}}$ inside some loop of $\Gamma$. Let $\mu(\cdot;\Upsilon_{k})$ denote the gasket measure on $\Upsilon_{k}$.
Proposition 2.6 ([MY26, Proposition 1.4]).
Let $\Gamma$ be a nested ${\rm CLE}_{\kappa^{\prime}}$ in a simply connected domain $D\subseteq\mathbf{C}$. Fix any $\ell\in\mathbf{N}$, $a>0$, and a compact set $K\subseteq D$. For each $r>0$, let $E_{r}$ be the event that
$\sup_{z\in K}\mu(B_{E}(z,r);\Upsilon_{k})\leq r^{d_{\rm CLE}-a}$
for each gasket $\Upsilon_{k}$ of $\Gamma$ of level at most $\ell$. Then
$\mathbf{P}[(E_{r})^{c}]=o^{\infty}(r)\quad\text{as }r\searrow 0.$
Proposition 2.7 ([MY26, Proposition 1.5]).
Let $\Gamma$ be a nested ${\rm CLE}_{\kappa^{\prime}}$ in a simply connected domain $D\subseteq\mathbf{C}$. Fix any $a>0$ and a compact set $K\subseteq D$. For each $r>0$, let $E_{r}$ be the event that the following holds. Let $\Upsilon_{k}$ be any of the gaskets of $\Gamma$, let $x\in\Upsilon_{k}\cap K$ and $w\in\partial B_{E}(x,r)\cap\partial B_{\mathrm{path}}(x,r)$. Let $B_{x,w}\subseteq\overline{B_{E}}(x,r)$ be a simply connected set containing all simple admissible paths within $\overline{B_{E}}(x,r)$ from $x$ to $w$. Then
$\mu(B_{x,w};\Upsilon_{k})\geq r^{d_{\rm CLE}+a}\quad\text{for each such $x,w$, and $k$.}$
Then
$\mathbf{P}[(E_{r})^{c}]=o^{\infty}(r)\quad\text{as }r\searrow 0.$
@@ 3.3.1. Approximation scheme
Let $\mu(\cdot;\Gamma)$ denote the gasket measure (see Section 2.3), and let $\mathfrak{V}_{\epsilon}$ be a Poisson point process with intensity measure $\lambda_{\epsilon}\mu(\cdot;\Gamma)$. As we will explain just below, the concentration results from [MY26] for $\mu(\cdot;\Gamma)$ imply that every Euclidean ball of radius $\epsilon$ is likely to contain many points of $\mathfrak{V}_{\epsilon}$ (i.e., at least a negative power of $\epsilon$).
@@ 4.3. Good annulus event
If we condition on the chords of a multichordal ${\rm CLE}_{\kappa^{\prime}}$, the remainder has the conditional law of an ordinary ${\rm CLE}_{\kappa^{\prime}}$. Hence, there exists some $\widetilde{D}$ with $\operatorname{dist_{E}}(z,\partial\widetilde{D})\in[2^{-j},2^{-j+1}]$ such that
@@ 5. Uniqueness
In the previous section, we proved that if we have two weak ${\rm CLE}_{\kappa^{\prime}}$ resistance forms $(\mathcal{E}^{V}(\cdot,\cdot;\Gamma))_{V\in\mathfrak{C}}$ and $(\widetilde{\mathcal{E}}^{V}(\cdot,\cdot;\Gamma))_{V\in\mathfrak{C}}$ with comparable scaling factors, then they are bi-Lipschitz equivalent with deterministic constants $0<c_{1}\leq c_{2}<\infty$. We show in Section 5.1 that in this setting one can take $c_{1}=c_{2}$. Then we show in Section 5.2 that there is a unique scaling exponent ${\alpha_{\mathrm{r}}}>0$ depending only on $\kappa^{\prime}$ such that every ${\rm CLE}_{\kappa^{\prime}}$ resistance form satisfies scaling covariance with this exponent ${\alpha_{\mathrm{r}}}>0$. This in particular lifts the assumption (4.1) on the comparability of the scaling factors.
@@ 5.2. Uniqueness of the exponent
In this section we show the scaling covariance of ${\rm CLE}_{\kappa^{\prime}}$ resistance forms with a scaling factor $\lambda^{{\alpha_{\mathrm{r}}}}$ where ${\alpha_{\mathrm{r}}}>0$ depends only on $\kappa^{\prime}$ (Proposition 5.5).
@@
There exists ${\alpha_{\mathrm{r}}}\in[{d_{{\mathrm{dbl}}}},{d_{{\rm SLE}}}]$ so that the following is true. Suppose that $(\mathcal{E}^{V}(\cdot,\cdot;\Gamma))_{V\in\mathfrak{C}}$ is a weak ${\rm CLE}_{\kappa^{\prime}}$ resistance form, then it is a (strong) ${\rm CLE}_{\kappa^{\prime}}$ resistance form with the scaling exponent $\alpha_{\mathrm{r}}$}.
We will prove Proposition 5.5 in two steps. The first step, stated and proved in Lemma 5.6 just below, is to show that each weak ${\rm CLE}_{\kappa^{\prime}}$ resistance form is a (strong) ${\rm CLE}_{\kappa^{\prime}}$ resistance form with some scaling exponent ${\alpha_{\mathrm{r}}}>0$. We will then prove that the value of ${\alpha_{\mathrm{r}}}>0$ does not depend on the choice of the ${\rm CLE}_{\kappa^{\prime}}$ resistance form, but only on the choice of $\kappa^{\prime}\in(4,8)$.
@@
Suppose that we have the setup of Section 4.1 but condition on the event that $\mathcal{L}\cap\partial\mathbf{D}=\emptyset$. [^7] This conditioning implies that for each $\lambda\in(0,1]$, the law of $\lambda\Gamma$ is absolutely continuous with respect to the law of $\Gamma$, see [MY25, Lemma 7.2]. Let $(\mathcal{E}^{V}(\cdot,\cdot;\Gamma))_{V\in\mathfrak{C}}$ be a weak ${\rm CLE}_{\kappa^{\prime}}$ resistance form. Then there exists ${\alpha_{\mathrm{r}}}\in[{d_{{\mathrm{dbl}}}},{d_{{\rm SLE}}}]$ such that the following holds. Let $\lambda\in(0,1]$ and define $(\widetilde{\mathcal{E}}^{V}(f,f;\Gamma))_{V\in\mathfrak{C}}=(\mathcal{E}^{\lambda V}(f(\lambda^{-1}\cdot),f(\lambda^{-1}\cdot);\lambda\Gamma))_{V\in\mathfrak{C}}$. Then almost surely $\widetilde{\mathcal{E}}^{V}(\cdot,\cdot;\Gamma)=\lambda^{-{\alpha_{\mathrm{r}}}}\mathcal{E}^{V}(\cdot,\cdot;\Gamma)$ for each $V\in\mathfrak{C}$.
@@
and therefore $c(\lambda^{\prime}\lambda)=c(\lambda^{\prime})c(\lambda)$. Moreover, $\lambda\mapsto c(\lambda)$ is strictly decreasing with $\lim_{\lambda\to 0}c(\lambda)=\infty$ due to Proposition 4.4. It follows that $c(\lambda)=\lambda^{-{\alpha_{\mathrm{r}}}}$ for some ${\alpha_{\mathrm{r}}}>0$. It also follows from Proposition 4.4 that ${\alpha_{\mathrm{r}}}\in[{d_{{\mathrm{dbl}}}},{d_{{\rm SLE}}}]$. ∎
@@
It remains to show that if $\mathcal{E}(\cdot,\cdot;\Gamma)$ and $\widetilde{\mathcal{E}}(\cdot,\cdot;\Gamma)$ are two weak ${\rm CLE}_{\kappa^{\prime}}$ resistance forms with respective scaling exponents ${\alpha_{\mathrm{r}}},\widetilde{{\alpha_{\mathrm{r}}}}$, then ${\alpha_{\mathrm{r}}}=\widetilde{{\alpha_{\mathrm{r}}}}$. Suppose that ${\alpha_{\mathrm{r}}}>\widetilde{{\alpha_{\mathrm{r}}}}$. Following the proof of Proposition 4.9, there is a covering by good regions $\{U_{l}\}$ analogously to Corollary 4.11 such that
$M^{-1}2^{j{\alpha_{\mathrm{r}}}}\leq\frac{\mathcal{E}^{U_{l}}(f,f;\Gamma)}{(\sup_{K_{l}}f-\inf_{K_{l}}f)^{2}}\leq M2^{j{\alpha_{\mathrm{r}}}}$
@@
$M^{-1}2^{j\widetilde{{\alpha_{\mathrm{r}}}}}\leq\frac{\widetilde{\mathcal{E}}^{U_{l}}(f,f;\Gamma)}{(\sup_{K_{l}}f-\inf_{K_{l}}f)^{2}}\leq M2^{j\widetilde{{\alpha_{\mathrm{r}}}}}$